# Recent Advances and Trends in Research Paper Recommender Systems: A Comprehensive Survey

### v1

@@ 3.4 What are the main challenges addressed by the RP-RSs?
Evaluation remains a significant challenge in RP-RSs due to the subjective nature of relevance, the absence of a definitive ground truth, and the difficulties associated with applying traditional evaluation metrics. Conventional metrics such as precision and recall fail to account for critical factors such as novelty, diversity, and user satisfaction. Furthermore, the cold start problem—where new users or papers lack sufficient interaction data—adds an additional layer of complexity to the evaluation process. As highlighted by [35], user evaluation is essential for assessing how well recommendations align with individual research needs, yet obtaining direct user feedback in academic contexts remains a challenge. Moreover, given the rapid evolution of research fields and the often indirect nature of feedback, evaluating how well these systems meet the dynamic demands of researchers becomes particularly difficult. Due to these challenges, there remains a scarcity of works that incorporate user-centered evaluation approaches. Recent works on RP-RSs involve user studies with varying configurations. [11] uses a three-point rating scale for paper similarity evaluation but provided no details about participants. [52] conducted a qualitative user study with four participants (two junior and two senior researchers) who evaluated publication similarity and complementarity across 20 starting points. [31] involved ten researchers in assessing the ability of the system to recommend older papers and those with no citations, but no further participant details were provided. [50] engaged 30 domain experts to evaluate recommendations across 10 research areas, with experts providing feedback on the relevance of the suggested articles. [15] conducted an online study with 112 participants, using a 1-5 like scale to assess recommendation quality, but no expert information was provided . [53] relied on experts from the Computer Science domain to manually rank articles for a chronological learning-based system, though no further expert details were given. [44] used a survey for participants to evaluate publication relevance, but no participant or expert details were provided.
@@
In previous surveys, it was highlighted that few approaches were available as prototypes, and there was often a disconnect between methods described in papers and real-world systems, a situation that remains unchanged to date. Among the works analyzed in this new survey, [33] emphasizes the scalability problem when applying state-of-the-art models. [33] states that while BERT models show outstanding performance in experiments using test data from model organism databases, the need for scalability led to the deployment of a simpler approach based on TF-IDF vectorization and a Naïve Bayes classifier. [57] points out that state-of-the-art machine learning RSs in the literature are typically limited to datasets with only a few thousand users and items, creating a significant gap between the datasets used in published systems and those available in real-world applications. [80] demonstrates the applicability of community detection to various scenarios in order to achieve more scalable results, which is crucial in the current age of digital libraries . Currently, only 9.52% of the works analyzed address this challenge by developing actual prototypes [50, 15, 52, 58, 33] or applying algorithms to real systems [44], while some others, plan to develop real web applications [54, 39]
@@
These issues are particularly evident in collaborative filtering systems, which rely on user interaction history to provide personalized suggestions. Recent advances have led to approaches that go beyond pure collaborative filtering, incorporating additional techniques to address these challenges. Current research follows the trends outlined in previous surveys. In recent research, 95.23% of works utilize content-based information to mitigate the cold start and sparsity issues. A significant portion of these works specifically targets overcoming the cold start problem [16, 43, 51, 75, 30, 46, 24, 17, 29, 13, 50] or the sparsity problem [60, 25]. Another set of works attempts to tackle both problems [74, 20, 56, 60, 57]. In some cases, test performance is carried out in cold start and sparse scenarios [77, 4, 20, 32, 57]
@@
The main privacy challenges in RP-RSs include protecting sensitive user data, ensuring proper anonymization without losing personalization, obtaining informed user consent, and securing data against breaches. Additionally, there is a lack of transparency in how user data is used, complicating privacy management. Compliance with privacy regulations, such as the General Data Protection Regulation (GDPR), adds further complexity, and there are concerns about potential biases and discrimination arising from how data is handled. The current situation remains largely the same, as many methods still overlook privacy concerns in their designs. No privacy conflicts were found in the works analyzed for this survey, as the data used in the various works are either public (see Table 3.2.1, Table 3.2.3, Table 3.2.2), labeled [59], anonymized (Table 3.2.4), or artificial [33, 57] .
@@
As stated in [35], numerous RP-RSs failed to clearly define their target audience, thus overlooking the distinct needs of various user groups in their evaluations. While some works identified specific audiences, such as junior or senior researchers, most did not adequately consider how these differences influenced the recommendation process. Recent works, however, provide some examples that address this issue. [21] introduces a student academic paper social recommendation model to address the interaction challenges between students and academic papers. [30] focuses on cold start junior researchers, recognizing the difficulty in learning research interests for users with limited publications. [76] targets postgraduates, [33] researches and curators, and [28] emphasize the need to differentiate between novice and experienced researchers in RP-RSs. [67] study researchers within scientific social networks, and [15] propose a ranking approach that distinguishes between new and experienced users. [55] focuses on recommending papers for researchers to cite. While these recent works reflect a growing awareness of the importance of target audience considerations, only about 12.69% of the analyzed works discuss or mention the target audience in their evaluations, highlighting that very few works address this aspect.
@@
[35] underlines the importance of explainability in RP-RSs   noting that while accuracy often takes priority, confidence and trust in the system are crucial for user satisfaction. Explainability enhances trust and engagement by making recommendations more transparent and understandable. However, in the works surveyed in this study, explainability remains largely unaddressed, and the increasing complexity of the proposed systems (see Section 3.1.3, Section 3.1.4, Section 3.1.5) further complicates the provision of clear and understandable justifications for recommendations.
@@
[35] highlights that RP-RSs often rely on a variety of datasets, many of which are proprietary and not publicly accessible. Although some approaches incorporate open datasets for evaluation, a considerable portion still fails to prioritize this aspect. They emphasize the importance of utilizing publicly available datasets or developing datasets that are shared and accessible to enhance the reproducibility of research in this domain. Even when public datasets are employed, modifications resulting from filtration or cleansing procedures—frequently lacking adequate documentation—compromise the ability to reproduce findings (see Table 3.2). Specifically, 9.90% of datasets are either modified or created by the authors and made publicly available, while 13.86% are accessible upon request. Furthermore, 27.72% of datasets, originally publicly available, are employed without any further modifications, and the remaining 47.52% consist of datasets that are either originally publicly available or ad-hoc constructed datasets, but are not accessible. These final datasets represent nearly 50% of all datasets employed, highlighting a significant barrier to reproducibility in the field.
@@
[35] discusses comparability issues in RP-RSs   noting that many approaches identify as RP-RSs but are often evaluated against general RSs or methodologies from different sub-domains. Emphasizing that even when approaches claim broader applicability, they should still be compared with systems specifically designed for paper recommendation to assess their relevance. Furthermore, cases where baselines are not established systems but artificial ones or entirely absent are highlighted. [35] stresses that evaluations of RP-RSs should always include at least one dedicated RP-RS to ensure a clear assessment of relevance within the field. Regarding the current situation, 28.57% of works utilize a combination of RP-RSs and general RSs baselines, followed by 25.40% that employ only RP-RSs baselines, and 17.46% that use RSs baselines. Notably, 15.87% of works do not compare recommendation algorithms but focus instead on paper and user representation algorithms. These trends still reflect ongoing challenges in ensuring proper comparability and comprehensive evaluation in the field.
@@ 3.5 What new challenges have arisen?
Although this issue was initially highlighted in the work of [9], it has not remained a primary focus in subsequent research. However, with the increasing interdisciplinarity of modern research, the need to accommodate the diversity of disciplines in RP-RSs has become an emerging challenge. Research on RP-RSs often relies on datasets that are limited to specific fields of study, which can hinder the generalization and applicability of these systems across various academic and research domains. Datasets containing papers from multiple disciplines are crucial for developing RP-RSs that can provide valuable suggestions to a broader audience. The research carried out indicates that only 28.91% of datasets are multidisciplinary, while 59.03% are not, and 12.04% do not provide information on the disciplines represented in the dataset. Furthermore, many works identify this issue as an aspect to be addressed in future research [76, 39, 30, 53, 20, 71]. This challenge should be reintroduced into the discussion, as it remains critical for the development of more generalizable and robust systems.
@@
Dynamic preferences refer to the evolving interests and needs of users over time. In the context of RP-RSs   academic interests of the researchers are subject to change, often influenced by factors such as new developments within their primary field, emerging topics, or the exploration of adjacent areas of study. As these preferences shift, RP-RSs must be capable of adapting to these changes in order to provide timely and relevant suggestions. Traditional RP-RSs   however, tend to assume a degree of stability in user preferences, which can limit their ability to reflect the dynamic and interdisciplinary nature of the academic journey of the researcher. To address this, RP-RSs need mechanisms that can track and predict the evolution of user preferences, ensuring recommendations align with the changing academic trajectories of users. In the current research, some works focus on incorporating dynamic trends in user interests by leveraging temporal data and heterogeneous information networks, while others explore the integration of attention mechanisms to capture evolving patterns in user behavior. Despite these efforts, only 15.87% of the works analyzed specifically tackle the issue of dynamic preferences [68, 69, 51, 30, 25, 49, 70, 55, 78, 15].
@@
The increasing availability of data from social media platforms offers significant opportunities to enhance RP-RSs. Social media data, such as the online activity of the researchers, collaborations, and discussions, can provide valuable insights into dynamic academic trends and shifting interests, complementing traditional sources such as citations and publication history. However, integrating such data presents challenges, including ensuring data quality, managing diverse and noisy information, and addressing privacy concerns. Recently, some works have begun exploring the use of social media and structural data in RP-RSs, focusing on group dynamics and collaboration networks to enhance recommendation accuracy and mitigate issues such as sparsity and cold-start problems. These works utilize techniques such as heterogeneous graph neural networks with hierarchical attention mechanisms [67], PMF with evidential reasoning [66], and hypergraph convolutional networks to model complex academic social structures [21]. The emergence of social media opens up new opportunities for research groups that are worth exploiting.
@@
Multitasking in RP-RSs refers to the simultaneous execution of multiple related tasks within a single model to enhance the accuracy and relevance of paper recommendations. This approach goes beyond simply recommending papers based on user preferences by integrating additional tasks. In the analyzed works, only 4 papers employed such characteristics: [32] proposes a multi-task RP-RS that simultaneously predicts paper recommendations and generates meta-data, such as keywords. Similarly, [45] combines models for academic paper link prediction and paper recommendation. Furthermore, [1] suggests the potential of extending their approach by integrating multi-task learning, where a single model is trained for both tag prediction and paper recommendation tasks. More recently, [71] introduces a graph neural network model to handle different tasks such as paper and co-author recommendation. While multitasking has been a concept scarcely mentioned in prior RP-RSs, it now appears to be gaining traction. This shift opens up new opportunities in the field, although it also introduces challenges, such as balancing task priorities, ensuring efficient resource use, and preventing interference between tasks.
What-if scenarios and synthetic data.
Handling what-if scenarios in RP-RSs is a key consideration for developing systems capable of predicting or providing recommendations based on hypothetical or alternative academic contexts. These scenarios may involve changes in the interests of the user, shifting collaboration networks, the emergence of new academic fields, or even the need for larger datasets to address scalability issues. For RP-RSs to effectively respond to these dynamic conditions, the datasets used need to reflect real-world academic environments that can simulate such changes. However, many existing datasets fail to capture this necessary complexity, limiting the ability of RP-RSs to adapt to evolving trends or shifting user needs. When algorithms or models are designed to address specific challenges in RP-RSs   it is crucial that the datasets used represent these specific scenarios. In the reviewed literature, synthetic data has not been employed, although its potential has been explored in some previous and peripheral works [10, 37, 2, 41, 12]. Exploring the full potential of synthetic data could provide new opportunities for developing more adaptable and scalable RP-RSs.

### v2

@@ 3.4 What are the main challenges addressed by the RP-RSs?
Evaluation remains a significant challenge in RP-RSs due to the subjective nature of relevance, the absence of a definitive ground truth, and the difficulties associated with applying traditional evaluation metrics. Conventional metrics such as precision and recall fail to account for critical factors such as novelty, diversity, and user satisfaction. Furthermore, the cold start problem—where new users or papers lack sufficient interaction data—adds an additional layer of complexity to the evaluation process. As highlighted by [IJ:DL:KREUTZ:2022], user evaluation is essential for assessing how well recommendations align with individual research needs, yet obtaining direct user feedback in academic contexts remains a challenge. Moreover, given the rapid evolution of research fields and the often indirect nature of feedback, evaluating how well these systems meet the dynamic demands of researchers becomes particularly difficult. Due to these challenges, there remains a scarcity of works that incorporate user-centered evaluation approaches. Recent works on RP-RSs involve user studies with varying configurations. [IJ:ELECTRONICS:Blazevic:2023] uses a three-point rating scale for paper similarity evaluation but provided no details about participants. [IC:WIST:Roßner:2023] conducted a qualitative user study with four participants (two junior and two senior researchers) who evaluated publication similarity and complementarity across 20 starting points. [IJ:Scientometrics:Kanwal:2024] involved ten researchers in assessing the ability of the system to recommend older papers and those with no citations, but no further participant details were provided. [IJ:IS:PINEDO:2024] engaged 30 domain experts to evaluate recommendations across 10 research areas, with experts providing feedback on the relevance of the suggested articles. [IJ:IS:Chaudhuri:2022] conducted an online study with 112 participants, using a 1-5 like scale to assess recommendation quality, but no expert information was provided . [IJ:IEEE:SARWAR:2021] relied on experts from the Computer Science domain to manually rank articles for a chronological learning-based system, though no further expert details were given. [J:JIS:MUNGEN:2022] used a survey for participants to evaluate publication relevance, but no participant or expert details were provided.
@@
In previous surveys, it was highlighted that few approaches were available as prototypes, and there was often a disconnect between methods described in papers and real-world systems, a situation that remains unchanged to date. Among the works analyzed in this new survey, [IJ:DATABASE:Kart:2022] emphasizes the scalability problem when applying state-of-the-art models. [IJ:DATABASE:Kart:2022] states that while BERT models show outstanding performance in experiments using test data from model organism databases, the need for scalability led to the deployment of a simpler approach based on TF-IDF vectorization and a Naïve Bayes classifier. [IJ:KIS:STERGIOPOULOS:2024] points out that state-of-the-art machine learning RSs in the literature are typically limited to datasets with only a few thousand users and items, creating a significant gap between the datasets used in published systems and those available in real-world applications. [IJ:JOI:Yadav:2022] demonstrates the applicability of community detection to various scenarios in order to achieve more scalable results, which is crucial in the current age of digital libraries . Currently, only 9.52% of the works analyzed address this challenge by developing actual prototypes [IJ:IS:PINEDO:2024, IJ:IS:Chaudhuri:2022, IC:WIST:Roßner:2023, IJ:IEEE:STERLING:2022, IJ:DATABASE:Kart:2022] or applying algorithms to real systems [J:JIS:MUNGEN:2022], while some others, plan to develop real web applications [ARXIV:Shen:2024, IJ:Heliyon:LI:2024]
@@
These issues are particularly evident in collaborative filtering systems, which rely on user interaction history to provide personalized suggestions. Recent advances have led to approaches that go beyond pure collaborative filtering, incorporating additional techniques to address these challenges. Current research follows the trends outlined in previous surveys. In recent research, 95.23% of works utilize content-based information to mitigate the cold start and sparsity issues. A significant portion of these works specifically targets overcoming the cold start problem [IJ:AS:Chen:2023, IJ:ESWA:MEI:2022, IJ:IEEE:REN:2022, IJ:SC:XI:2024, IJ:Scientometrics:Jiang:2023, IJ:SECS:Hamisu:2024, IJ:JS:Huang:2023, IJ:AI:GAO:2023, IJ:AUT:Jafari:2023, IJ:EIT:Smail:2023, IJ:IS:PINEDO:2024] or the sparsity problem [IC:ICITACEE:Switrayana:2022, J:EIJ:Huang:2024]. Another set of works attempts to tackle both problems [IJ:ESWA:XI:2023, IJ:TIIS:GUO:2024, IJ:DSM:Sivasankari:2023, IC:ICITACEE:Switrayana:2022, IJ:KIS:STERGIOPOULOS:2024]. In some cases, test performance is carried out in cold start and sparse scenarios [IJ:ESWA:XIAO:2023, IJ:Scientometrics:ALI:2022, IJ:TIIS:GUO:2024, arXiv:KARIMI:2022, IJ:KIS:STERGIOPOULOS:2024]
@@
The main privacy challenges in RP-RSs include protecting sensitive user data, ensuring proper anonymization without losing personalization, obtaining informed user consent, and securing data against breaches. Additionally, there is a lack of transparency in how user data is used, complicating privacy management. Compliance with privacy regulations, such as the General Data Protection Regulation (GDPR), adds further complexity, and there are concerns about potential biases and discrimination arising from how data is handled. The current situation remains largely the same, as many methods still overlook privacy concerns in their designs. No privacy conflicts were found in the works analyzed for this survey, as the data used in the various works are either public (see Table 3.2.1, Table 3.2.3, Table 3.2.2), labeled [JCDL:Sugiyama:2013], anonymized (Table 3.2.4), or artificial [IJ:DATABASE:Kart:2022, IJ:KIS:STERGIOPOULOS:2024] .
@@
As stated in [IJ:DL:KREUTZ:2022], numerous RP-RSs failed to clearly define their target audience, thus overlooking the distinct needs of various user groups in their evaluations. While some works identified specific audiences, such as junior or senior researchers, most did not adequately consider how these differences influenced the recommendation process. Recent works, however, provide some examples that address this issue. [IJ:mathematics:GUO:2024] introduces a student academic paper social recommendation model to address the interaction challenges between students and academic papers. [IJ:Scientometrics:Jiang:2023] focuses on cold start junior researchers, recognizing the difficulty in learning research interests for users with limited publications. [IJ:AI:Xiao:2023] targets postgraduates, [IJ:DATABASE:Kart:2022] researches and curators, and [IC:IKT:JAFARI:2021] emphasize the need to differentiate between novice and experienced researchers in RP-RSs. [IJ:ASC:Wang:2024] study researchers within scientific social networks, and [IJ:IS:Chaudhuri:2022] propose a ranking approach that distinguishes between new and experienced users. [IC:ASSP:Shen:2024] focuses on recommending papers for researchers to cite. While these recent works reflect a growing awareness of the importance of target audience considerations, only about 12.69% of the analyzed works discuss or mention the target audience in their evaluations, highlighting that very few works address this aspect.
@@
[IJ:DL:KREUTZ:2022] underlines the importance of explainability in RP-RSs   noting that while accuracy often takes priority, confidence and trust in the system are crucial for user satisfaction. Explainability enhances trust and engagement by making recommendations more transparent and understandable. However, in the works surveyed in this study, explainability remains largely unaddressed, and the increasing complexity of the proposed systems (see Section 3.1.3, Section 3.1.4, Section 3.1.5) further complicates the provision of clear and understandable justifications for recommendations.
@@
[IJ:DL:KREUTZ:2022] highlights that RP-RSs often rely on a variety of datasets, many of which are proprietary and not publicly accessible. Although some approaches incorporate open datasets for evaluation, a considerable portion still fails to prioritize this aspect. They emphasize the importance of utilizing publicly available datasets or developing datasets that are shared and accessible to enhance the reproducibility of research in this domain. Even when public datasets are employed, modifications resulting from filtration or cleansing procedures—frequently lacking adequate documentation—compromise the ability to reproduce findings (see Table 3.2). Specifically, 9.90% of datasets are either modified or created by the authors and made publicly available, while 13.86% are accessible upon request. Furthermore, 27.72% of datasets, originally publicly available, are employed without any further modifications, and the remaining 47.52% consist of datasets that are either originally publicly available or ad-hoc constructed datasets, but are not accessible. These final datasets represent nearly 50% of all datasets employed, highlighting a significant barrier to reproducibility in the field.
@@
[IJ:DL:KREUTZ:2022] discusses comparability issues in RP-RSs   noting that many approaches identify as RP-RSs but are often evaluated against general RSs or methodologies from different sub-domains. Emphasizing that even when approaches claim broader applicability, they should still be compared with systems specifically designed for paper recommendation to assess their relevance. Furthermore, cases where baselines are not established systems but artificial ones or entirely absent are highlighted. [IJ:DL:KREUTZ:2022] stresses that evaluations of RP-RSs should always include at least one dedicated RP-RS to ensure a clear assessment of relevance within the field. Regarding the current situation, 28.57% of works utilize a combination of RP-RSs and general RSs baselines, followed by 25.40% that employ only RP-RSs baselines, and 17.46% that use RSs baselines. Notably, 15.87% of works do not compare recommendation algorithms but focus instead on paper and user representation algorithms. These trends still reflect ongoing challenges in ensuring proper comparability and comprehensive evaluation in the field.
@@ 3.5 What new challenges have arisen?
Although this issue was initially highlighted in the work of [IJ:DL:BEEL:2015], it has not remained a primary focus in subsequent research. However, with the increasing interdisciplinarity of modern research, the need to accommodate the diversity of disciplines in RP-RSs has become an emerging challenge. Research on RP-RSs often relies on datasets that are limited to specific fields of study, which can hinder the generalization and applicability of these systems across various academic and research domains. Datasets containing papers from multiple disciplines are crucial for developing RP-RSs that can provide valuable suggestions to a broader audience. The research carried out indicates that only 28.91% of datasets are multidisciplinary, while 59.03% are not, and 12.04% do not provide information on the disciplines represented in the dataset. Furthermore, many works identify this issue as an aspect to be addressed in future research [IJ:AI:Xiao:2023, IJ:Heliyon:LI:2024, IJ:Scientometrics:Jiang:2023, IJ:IEEE:SARWAR:2021, IJ:TIIS:GUO:2024, IC:SEAI:Wang:2024]. This challenge should be reintroduced into the discussion, as it remains critical for the development of more generalizable and robust systems.
@@
Dynamic preferences refer to the evolving interests and needs of users over time. In the context of RP-RSs   academic interests of the researchers are subject to change, often influenced by factors such as new developments within their primary field, emerging topics, or the exploration of adjacent areas of study. As these preferences shift, RP-RSs must be capable of adapting to these changes in order to provide timely and relevant suggestions. Traditional RP-RSs   however, tend to assume a degree of stability in user preferences, which can limit their ability to reflect the dynamic and interdisciplinary nature of the academic journey of the researcher. To address this, RP-RSs need mechanisms that can track and predict the evolution of user preferences, ensuring recommendations align with the changing academic trajectories of users. In the current research, some works focus on incorporating dynamic trends in user interests by leveraging temporal data and heterogeneous information networks, while others explore the integration of attention mechanisms to capture evolving patterns in user behavior. Despite these efforts, only 15.87% of the works analyzed specifically tackle the issue of dynamic preferences [IC:DASFAA:Wang:2022, IJ:ESWA:Wang:2023, IJ:IEEE:REN:2022, IJ:Scientometrics:Jiang:2023, J:EIJ:Huang:2024, AC:ACIIDS:PAN:2024, IJ:AS:Wang:2024, IC:ASSP:Shen:2024, IJ:KBS:Xiao:2023, IJ:IS:Chaudhuri:2022].
@@
The increasing availability of data from social media platforms offers significant opportunities to enhance RP-RSs. Social media data, such as the online activity of the researchers, collaborations, and discussions, can provide valuable insights into dynamic academic trends and shifting interests, complementing traditional sources such as citations and publication history. However, integrating such data presents challenges, including ensuring data quality, managing diverse and noisy information, and addressing privacy concerns. Recently, some works have begun exploring the use of social media and structural data in RP-RSs, focusing on group dynamics and collaboration networks to enhance recommendation accuracy and mitigate issues such as sparsity and cold-start problems. These works utilize techniques such as heterogeneous graph neural networks with hierarchical attention mechanisms [IJ:ASC:Wang:2024], PMF with evidential reasoning [IJ:SMC:WANG:2022], and hypergraph convolutional networks to model complex academic social structures [IJ:mathematics:GUO:2024]. The emergence of social media opens up new opportunities for research groups that are worth exploiting.
@@
Multitasking in RP-RSs refers to the simultaneous execution of multiple related tasks within a single model to enhance the accuracy and relevance of paper recommendations. This approach goes beyond simply recommending papers based on user preferences by integrating additional tasks. In the analyzed works, only 4 papers employed such characteristics: [arXiv:KARIMI:2022] proposes a multi-task RP-RS that simultaneously predicts paper recommendations and generates meta-data, such as keywords. Similarly, [IC:EITCE:NIU:2023] combines models for academic paper link prediction and paper recommendation. Furthermore, [ARXIV:Mohamed:2022] suggests the potential of extending their approach by integrating multi-task learning, where a single model is trained for both tag prediction and paper recommendation tasks. More recently, [IC:SEAI:Wang:2024] introduces a graph neural network model to handle different tasks such as paper and co-author recommendation. While multitasking has been a concept scarcely mentioned in prior RP-RSs, it now appears to be gaining traction. This shift opens up new opportunities in the field, although it also introduces challenges, such as balancing task priorities, ensuring efficient resource use, and preventing interference between tasks.
Handling what-if scenarios in RP-RSs is a key consideration for the development of systems capable of predicting or providing recommendations under hypothetical or alternative academic contexts. These scenarios may involve changes in user interests, shifts in collaboration networks, the emergence of new academic fields, or the need for larger datasets to address scalability challenges. For RP-RSs to effectively respond to such dynamic conditions, the datasets employed must reflect realistic academic environments that are able to simulate these variations. However, many existing datasets lack the necessary complexity to capture these dynamics, which limits the adaptability of RP-RSs to evolving trends and changing user needs. When algorithms or models are designed to tackle specific challenges within RP-RSs, it is therefore essential that the datasets used accurately represent the corresponding scenarios. Despite its relevance for modeling controlled or hypothetical academic settings, synthetic data has not been explicitly considered in the reviewed works. Exploring its full potential may enable the development of more adaptable and scalable RP-RSs.